OpenStudy (anonymous):

\[\frac{ x }{ x ^{2}-3x-4 }+\frac{ 3 }{ x-4 }\]

OpenStudy (anonymous):

x^2 -3x - 4 = x^2 -4x+x -4 = x(x-4)+1(x-4) = (x+1)(x-4)

OpenStudy (anonymous):

does this help ?

OpenStudy (anonymous):

\[\frac{ 3x+4 }{ x ^{2}-4x-3 }\]

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

do it again

OpenStudy (anonymous):

and you changed the denominator for some reason :|

OpenStudy (anonymous):

\[\frac{ 4x+3 }{ x ^{2}-3x-4 }\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

\[\frac{ 4x-5 }{ 12x+4 \frac{ 3x-1 }{ 3x+1 } }\]

OpenStudy (anonymous):

have you tried?

OpenStudy (anonymous):

\[\frac{ x-16 }{ 4\left( 3x+1 \right) } ;x \neq -\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

lets do this together

OpenStudy (anonymous):

\[\frac{ (4x-5)(3x+1) }{ 12x(3x+1) + 4(3x-1) }\]

OpenStudy (anonymous):

that's the first step you did as well ?

OpenStudy (anonymous):

in fact there is nothing much to simplify here .. are you sure that you copied the question correct ?